OpenStudy (anonymous):

What is the solution to the rational equation x/5 + 2/5 = 7?

OpenStudy (maheshmeghwal9):

\[\frac{x}{5}+\frac{2}{5}=7.\]\[\implies \frac{x+2}{5}=7.\]\[\implies (x+2)=35.\] So; \[\implies \color{blue}{x= \space ?}\]

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

You need to move 2/5 to the other side, to do so we subtract it from both sides like so: \[x/5 + 2/5 - 2/5 = 7 - 2/5\]This simplifies to \[x/5 = 7 - 2/5\]and to simplify the right side, we make both terms have the same denominator by multiplying seven by five like so:\[x/5 = 35/5 - 2/5\] which becomes \[x/5 = 33/5\] and lastly, we multiply both sides by five: \[x = 33\]

OpenStudy (anonymous):

x = 33 i never thought of that

OpenStudy (maheshmeghwal9):

yeah

OpenStudy (anonymous):

what i did was: 2x/5 = 7 lol

OpenStudy (anonymous):

thanks

OpenStudy (maheshmeghwal9):

yw:)